OpenStudy (ilovebmth1234):

In the following reaction, how many grams of ferrous sulfide (FeS) will produce 0.56 grams of iron (III) oxide (Fe2O3)? 4FeS + 7O2 d 2Fe2O3 + 4SO2 The molar mass of ferrous sulfide is 87.92 grams and that of iron (III) oxide is 159.7 grams.

OpenStudy (ilovebmth1234):

A. 0.1541 grams B. 0.6166 grams C. 0.5085 grams D. 2.0344 grams

OpenStudy (ilovebmth1234):

@Jack_Prism

OpenStudy (rushwr):

Show me ur work and ur answer !

OpenStudy (ilovebmth1234):

@magicremix123

OpenStudy (rushwr):

look at the mole ratio between Fe2O3 and FeS that is 2:4 right? We can find the no of moles of Fe2O3 moles= mass divided by molar mass Moles of Fe2)3 = o.56g/ 159.7= 0.003506moles As we can see FeS has twice the moles as Fe2O3 so moles of FeS = 0.00351*2 = 0.007013 Moles = mass/molar mass mass= moles*molar mass mass of FeS = 0.007013 * 87.92= 0.6166g So the answer is B

OpenStudy (anonymous):

I would say c.) Thats what i got .....mhm

OpenStudy (anonymous):

Maybe, i did something wrong in the problem